What is the slope of the line that is parallel to 3x - 4y = -40?
Answer by TimothyLamb(4379)

3x - 4y = -40
4y = 3x + 40
y = (3/4)x + 10
---
answer:
slope = 3/4
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---